### v3

# Elias’ Encoding from Lagrangians and Renormalization

Abstract
In the present paper we give a principled derivation of Elias’ Omega code by combining a constrained variational formulation of prefix coding with a renormalization flow on codeword distributions. Starting from a Lagrangian that minimizes average code length under the Kraft–McMillan constraint, we show that the implied distribution is a fixed point of a coarse–graining map, yielding the canonical iterated logarithm length, up to an additive constant.
This establishes completeness and asymptotic optimality, and connects universal integer coding with coarse–grained entropy, uncertainty–type bounds, and multiplicity relations familiar from statistical physics. The renormalization operator induces a discrete flow that converges to the Elias fixed point for any admissible initialization, up to a bounded error, offering a clean bridge between information–theoretic constraints and RG–style scale invariance.
1 Introduction
Among several possible encoding procedures for integers, in his paper [1] Elias introduces the so–called $\omega$ encoding. This is a universal prefix code obtained by the simple iterative procedure below. Let $s=\text{bin}(x)$ be the binary string representation of a decimal number $x$, and let $x=\text{dec}(s)$ be the decimal value corresponding to a binary string $s$.
Algorithm 1 Elias’ Omega: Encoding
1:A positive integer $N\geq 1$
2:Encoded binary string for NN
3:Initialize an empty code string $C=\text{`'}$
4:Append a "0" to the end of $C$: $C\leftarrow C+\text{`0'}$ $\triangleright$ Termination marker
5:while $N>1$ do
6:  Prepend the binary representation of NN to $C$: $C\leftarrow\text{bin}(N)+C$
7:  Set $N\leftarrow\text{length}(\text{bin}(N))-1$ $\triangleright$ Number of bits in $\text{bin}(N)$, minus one
8:end while
9:return $C$
Algorithm 2 Elias’ Omega: Decoding
1:Encoded binary string $S$
2:Decoded positive integer NN
3:Initialize $N\leftarrow 1$
4:Initialize a pointer $i\leftarrow 1$ (to track the current position in $S$)
5:while $S[i]\neq\text{`0'}$ do
6:  Read $N^{\prime}\leftarrow\text{dec}(S[i:i+N])$
7:  Move the pointer $i\leftarrow i+N+1$
8:  Rewrite $N\leftarrow N^{\prime}$
9:end while
10:return NN
Since Shannon’s foundational work [2] and the Kraft–McMillan inequality [3, 4], a central question has been how to construct codes that approach entropy without knowledge of the true source distribution: such codes are called universal. Elias’s $\gamma$, $\delta$, and $\omega$ codes are all universal [1].
In this note, we show that $\omega$ follows directly from a constrained Lagrangian optimization combined with a scale‐invariance requirement for the implied distribution, yielding a clean derivation from fundamental physical and code-theoretic principles.
2 Optimal codes from Lagrangians and Renormalization
2.1 Lagrangian Optimization for Code Length
Some interesting and practically common statistical laws appear to follow as optimal solutions of Lagrangians combining together the Minimal Entropy Principle with various combinatorial or physical constraints. One notable example is Visser’s derivation of Zipf’s law in [5]. Another one is Kafri’s work of Benford’s law [6] (see also [7]).
Let us consider codes as probability spaces of codewords $\mathcal{C}\subseteq[1,+\infty)$, with the source distribution probability $p(x)$ of a codeword $x$ being a generalized function
$p(dx)\;=\;\sum_{w\in W}P(w)\,\delta_{w}(dx)\;+\;\rho(x)dx,$ (1)
i.e. a weighted sum of delta-functions together with a continuous probability density. Here $W$ is at most countable, $\rho\geq 0$ is integrable, and
$\sum_{w\in W}P(w)+\int_{\mathcal{C}}\rho(x)\,dx=1.$ (2)
Moreover, $W\cap\mathrm{supp}\,\rho=\emptyset$.
Fix an output alphabet $\Sigma_{D}$ of size $D\geq 2$, and let $\Sigma_{D}^{*}$ be its Kleene closure. The $D$-ary code length $\ell_{D}(x)$ can be replaced by the baseless (or nat) code length $\ell(x)=\ell_{D}(x)/\log D$, which we will prefer in the sequel.
Then the average code length equals
$\mathbb{E}[\ell]=\int_{\mathcal{C}}\ell(x)p(dx).$ (3)
A prefix code (or a prefix-free code) is a countable set $U\subset\Sigma_{D}^{*}$ of finite strings over a $D$-ary alphabet such that no $u\in U$ is a prefix of another $v\in U$.
Let us formally recast the Kraft–McMillan inequality [3, 4] for prefix codes in the continuous form:
$\int_{\mathcal{C}}\exp(-\ell(x))\nu(dx)\leq 1,$ (4)
where
$\nu(dx)=\sum_{w\in W}\delta_{w}(dx)+dx,$ (5)
is the Kraft measure.
Here, if $x\notin\mathcal{C}$ is not a codeword, we assume $\ell(x)=+\infty$. Let us note that for countable codes $p(x)$ is a linear combination of delta-functions, and the integrals above become the usual discrete sums. More details on prefix-free codes with mixed source distributions are provided in Appendix 3.
We shall add this condition as a constraint to the Lagrangian below:
$\mathcal{L}(\ell)=\int_{\mathcal{C}}\ell(x)p(dx)-\lambda\left(1-\int_{\mathcal{C}}\exp(-\ell(x))\nu(dx)\right)\to\min_{\ell},$ (6)
with $\lambda>0$. This implies that violating the Kraft–McMillan inequality would imply a penalty on the Lagrangian.
The whole Lagrangian will then minimize the average code length under the prefix-free constraint. This entropy–optimality interpretation is classical [8, 9, 10].
Since $p\ll\nu$, define the Radon–Nikodym derivative
$r(x)=\frac{dp}{d\nu}(x)=\begin{cases}P(w),&x=w\in W,\\[4.0pt] \rho(x),&x\in\mathrm{supp}\,\rho.\end{cases}$ (7)
Taking the functional derivative gives
$\frac{\delta\mathcal{L}}{\delta\ell}=r(x)-\lambda\exp(-\ell(x))=0\;(\nu\text{-a.e.}),$ (8)
while
$1=p(\mathcal{C})=\int_{\mathcal{C}}r(x)\nu(dx)=\lambda\int_{\mathcal{C}}\exp(\ell(x))\nu(dx)=\lambda.$ (9)
Indeed, the complementary slackness condition implies that the code so derived has to saturate Kraft’s inequality, and thus be complete (see [11] for fundamentals of convex analysis).
This implies $p(x)$ being the code’s implied probability with
$\ell(x)=\begin{cases}-\log P(w),&x=w\in W,\\ -\log\rho(x),&x\in\mathrm{supp}\,\rho.\end{cases}$ (10)
As a result, we obtain the following theorem.
Theorem 2.1.
Let $\mathcal{C}$ be a self-delimiting code with source distribution probability $p(x)$ and codeword length $\ell(x)$. If $\ell(x)$ is chosen to minimize the average code length, then the code is complete and the codeword probability is exactly the implied one.
2.2 Renormalizing the Codeword Distribution
Let us define the following renormalization operator $\mathcal{R}$ for probabilities that is consistent with exponentiation. Namely, we want to be able to renormalize for $\exp(x)$ if we know the probabilities for $x$. This, on the surface, does not imply the universality property shown in [1], but rather guarantees that the conditional distribution is the same over all orders of magnitude. However, we shall see that universality will readily follow from this simple distributional condition.
The process used below is in essence the one of “coarse‐graining” by replacing $x$ with $\log x$ followed by “renormalizing” the resulting probability (or code length). However, it appears to be technically much simpler than the classical renormalization techniques in physics [12, 13, 14] due to its code-theoretical nature.
Thus, put
$\mathcal{R}p(\exp(x))=p(x)\cdot\exp(-x),$ (11)
with the obvious fixed point given by a distribution $p(x)$ satisfying
$p(\exp(x))=p(x)\cdot\exp(-x).$ (12)
Note that the above can also be rewritten as a statement about the code length instead:
$\ell(\exp(x))=\ell(x)+x.$ (13)
The above, together with the natural constraint $\ell(x)\geq 0$, implies that $\ell(x)$ is given by the sum of iterated logarithms $\log(\log(\ldots\log(x)\ldots)$ that continues until we obtain a negative number.
This implies that
$\ell(x)=\log x+o(\log x),\text{ as }x\to\infty,$ (14)
and, moreover, we have that for any $\varepsilon>0$ and all $x\in\mathcal{C}$ large enough $o(\log x)\leq\varepsilon\log x$ and thus
$\int_{\mathcal{C}}o(\log x)p(dx)\leq O(1)+\varepsilon\int_{\mathcal{C}}\log(x)p(dx)\leq O(1)+\varepsilon H(p)=O(1),$ (15)
as $\varepsilon>0$ can be chosen arbitrarily.
This in its turn yields
$\mathbb{E}[\ell]=\int_{\mathcal{C}}\ell(x)\,p(dx)=\int_{\mathcal{C}}\log(x)\,p(dx)+O(1)\leq H(p)+O(1),$ (16)
where the last step is Wyner’s inequality [15]. Here,
$H(p)=-\int_{\mathcal{C}}\log p(x)\,p(dx)$ (17)
is the Shannon entropy [2] of $\mathcal{C}$. On the other hand, Shannon’s classical lower bound $\mathbb{E}[\ell]\geq H(p)$ also holds (Appendix 3.5).
Thus, we obtain the following statement.
Theorem 2.2.
Let $\mathcal{C}$ be a code for which the implied probability $p(x)$ of a codeword $x\in\mathcal{C}$ satisfies the renormalization condition $p(\exp(x))=p(x)\cdot\exp(-x)$. Then the code length of $\mathcal{C}$ behaves as $\ell(x)=\log x+o(\log x)$, for $x\gg 1$, and $\mathcal{C}$ is asymptotically optimal: $H(p)\leq\mathbb{E}[\ell]\leq H(p)+O(1)$.
We shall call any code satisfying Theorems 2.1 - 2.2 an Elias’ Omega code, for short.
2.3 Renormalization Flow Universality
Let the renormalization operator $\mathcal{R}$ act on an arbitrary codeword distribution $p(x)$ as
$\mathcal{R}[p](\exp(x))\;=\;p(x)\,\exp(-x),$ (18)
where the map $\log:\exp(x)\to x$ implements a coarse‐graining step or scale reduction, while the factor $\exp(-x)$ ensures $\int_{0}^{\infty}\mathcal{R}[p](x)\,dx=1$.
Assuming that $p(x)$ is implied, we have the induced operator for the code length $\ell(x)$ is
$T[\ell](x)\;=\;\ell\bigl(\log x\bigr)\;+\;\log x.$ (19)
Starting from any given code length function $\ell_{0}(x)$, we can transform it by setting
$\ell_{k+1}(x)=T[\ell_{k}](x)=\ell_{k}(\log x)+\log x,$ (20)
for $k=0,1,2,\ldots$, which gives us a discrete flow.
The fixed point $\ell^{*}(x)$ of this flow is exactly the Elias’ code length discussed above:
$\ell_{*}(x)=T[\ell_{*}](x)=\ell_{*}(\log x)+\log x.$ (21)
Recursively unrolling this relation yields exactly
$\ell_{*}(x)=\log x\;+\;\log(\log x)\;+\;\log\bigl(\log(\log x)\bigr)\;+\;\cdots,$ (22)
terminating when the next $\log$ argument drops below $1$, i.e. the “0” termination of Elias’ Omega encoding.
Since $\log x\leq x/2$ for any $x\geq 0$, we get that
$\ell_{*}(x)\leq\log x\cdot\sum^{m}_{i=0}\frac{1}{2^{i}}\leq 2\log x$ (23)
where $m=\log^{*}x$, the iterated logarithm of $x$, known for its very low rate of growth. This implies that
$\ell_{*}(x)=O(\log x),$ (24)
all the while being a self-delimiting code unlike the usual binary encoding, which obviously also has $O(\log x)$ code length.
Moreover, given a starting non-negative function $\ell_{0}(x)$ on the codewords of $\mathcal{C}$, repeated application of the map
$T:\ell\mapsto\ell\circ\log+\log$ (25)
flows to the point
$\ell_{m}(x)=\sum_{i=0}^{m-1}x_{i}+\ell_{0}(x_{m}),$ (26)
where
$x_{i+1}$ $=\log x_{i},\text{ for }i=0,\dots,m-1,$
$\text{with }x_{0}=x\text{ and }x_{m}\leq 1.$
Thus
$|\ell_{*}(x)-\ell_{m}(x)|\leq\ell_{0}(x_{m})=O(1),$ (27)
since $\ell_{0}(x_{m})$ is simply the code length of the termination marker.
We can summarize our findings as the following propositions.
Proposition 2.3.
Let $\mathcal{C}$ be an Elias’ Omega code with code length $\ell(x)$ for $x\in\mathcal{C}$. Then
$\ell(x)=\log x+\log^{(2)}x+\ldots+\log^{(m)}x,$
with $m=\log^{*}x$, the iterated logarithm of $x$.
Proposition 2.4.
Let $\ell_{0}(x)$ be some non-negative function on the codewords of an Elias’ Omega code $\mathcal{C}$. Then the repeated iterations of the renormalization map $\ell\mapsto\ell\circ\log+\log$ almost converge to the canonical code length. Namely, if $m=\log^{*}x$, let $\ell_{m}=T^{(m)}[\ell]$, and then
$|\ell(x)-\ell_{m}(x)|<O(1)$
independent of $x$.
2.4 Passing to the Discrete Encoding
Naturally, if we now recast our observation into the discrete setting, Elias’ Omega encoding [1] will provide a practical realization with very close properties.
As it was mentioned before, Elias’ Omega encoding has to be complete. Indeed, let $L(n)$ be the length of $n$ in Elias’ Omega encoding, and let $\beta(n)=\lfloor\log_{2}n\rfloor+1$ be the binary length of $n$. Then we have
$L(n)=L(\beta(n)-1)+\beta(n).$ (28)
Let also
$I_{k}=\{n\in\mathbb{N}\,|\,\beta(n)=k\}=\{n\in\mathbb{N}\,|\,2^{k-1}\leq n\leq 2^{k}-1\}.$ (29)
We have that $I_{k}$ contains $2^{k-1}$ elements, and thus
$S_{k}=\sum_{n\in I_{k}}2^{-L(n)}=\sum_{n\in I_{k}}2^{-L(k-1)-k}=2^{-L(k-1)-1}.$ (30)
Then for the Kraft’s sum $S=\sum_{n\in\mathbb{N}}2^{-L(n)}$ we get
$S=\sum^{\infty}_{k=1}S_{k}=S_{1}+\sum^{\infty}_{k=2}S_{k}=\frac{1}{2}+\sum^{\infty}_{k=2}2^{-L(k-1)-1}=\frac{1}{2}+\frac{1}{2}\sum^{\infty}_{k=1}2^{-L(k)}=\frac{1+S}{2}.$ (31)
This readily implies that $S=1$. The real rate of convergence is very slow though: in our computations we have reached (see [16])
$\sum^{2^{2^{24}}}_{n=1}2^{-L(n)}\approx 0.9697265625\ldots$ (32)
Acknowledgments
This material is based upon work supported by the Google Cloud Research Award number GCP19980904.
3 Appendix: Kraft–McMillan and Quantization
3.1 Mixed Source Distributions
Let $p(x)$ be a probability measure on the set of codewords $\mathcal{C}=[1,+\infty)$ of the form
$p(x)\;=\;\sum_{w\in W}P(w)\,\delta_{w}(x)\;+\;\rho(x),$
where $W$ is at most countable, $\rho\geq 0$ is integrable, $W\cap\mathrm{supp}\,\rho=\emptyset$, and
$\sum_{w\in W}P(w)+\int_{\mathcal{C}}\rho(x)\,dx=1.$
Fix an output alphabet $\Sigma_{D}$ of size $D\geq 2$, and let $\Sigma_{D}^{*}$ be its Kleene closure. A prefix code (or a prefix-free code) is a countable set $U\subset\Sigma_{D}^{*}$ of finite strings over a $D$-ary alphabet such that no $u\in U$ is a prefix of another $v\in U$.
Instead of the $D$-ary code length $\ell_{D}(x)$ for $x\in\mathcal{C}$, we shall use the baseless (more precisely, base–invariant) code length $\ell(x)=\ell_{D}(x)/\log D$.
In the mixed (discrete and continuous) distribution case, we shall assign source outcomes to these finite strings by grouping outcomes into finitely many coding cells represented either by singleton atoms or quantization cells of the density. Then the notion of prefix-freeness will carry over from the standard one on the finite codebook $U$.
We say that $\mathcal{C}$ satisfies the Kraft–McMillan condition if
$\int_{\mathcal{C}}\exp(-\ell(x))dx\leq 1,$
where $\ell(x)$ is the code length of $x\in\mathcal{C}$.
3.2 Continuous Code Quantization
Let us choose
  • (i)
    a finite subset $W_{0}\subset W$, and
  • (ii)
    a finite measurable partition $\Pi=\{A_{1},\dots,A_{m}\}$ of some measurable $A\subseteq\{x:\rho(x)>0\}$.
Then $W_{0}$ is of atoms to encode explicitly, and $\Pi$ is a quantization of the continuous support.
Next, let us define the finite index set
$S\;=\;W_{0}\,\cup\,\{A_{1},\dots,A_{m}\},$
and ascribe the code lengths to its elements
$\ell_{s}\;=\;\begin{cases}\ell(w),&s=w\in W_{0},\\[2.0pt] \displaystyle-\log\int_{A_{j}}\exp(-\ell(x))\,dx,&s=A_{j}\in\Pi.\end{cases}$
3.3 Prefix-free Quantized Codes
A $D$-ary prefix-free quantized code for the mixed law restricted to $S$ is a pair $(U,\{S_{u}\}_{u\in U})$ where:
  • $U\subset\Sigma_{D}^{*}$ is finite and prefix-free;
  • $\{S_{u}\}_{u\in U}$ is a measurable partition of the set represented by $S$, i.e. atoms in $W_{0}$ and cells in $\Pi$, with $\pi_{u}=p(S_{u})>0$ for all $u\in U$.
As before, we shall use base-invariant notation or, equivalently, base $D=e$ arity.
Theorem. With $S$, $\{\pi_{s}\}$ and $\{\ell_{s}\}$ as above, we have
$\sum_{s\in S}\exp(-\ell_{s})\;\leq\;1,$
if $\mathcal{C}$ satisfies the Kraft–McMillan condition.
Indeed, by definition,
$\sum_{s\in S}\exp(-\ell_{s})=\sum_{W_{0}}\exp(-\ell(w))+\sum_{\Pi}\int_{A_{j}}\exp(-\ell(x))dx\leq$
$\leq\sum_{W_{0}}\exp(-\ell(w))+\int_{\Pi}\exp(-\ell(x))dx=\int_{\mathcal{C}}\exp(-\ell(x))dx\leq 1.$
Thus, there exists a quantized prefix code for the mixed-distribution code $\mathcal{C}$, for every admissible choice of the index set $S$ as above.
3.4 Heisenberg–style and Boltzmann–style relations
Let the Kraft sum be $K=\sum_{s}e^{-\ell_{s}}$, and let us define the Kraft distribution as
$q_{s}=e^{-\ell_{s}}/K,$
which is a code-length-related Gibbs measure.
Then we have the exact identity
$\bar{\ell}=\sum_{s\in S}\pi_{s}\ell_{s}\;=\;H(\pi)\;+\;D_{\mathrm{KL}}(\pi\|q)\;+\;\ln\frac{1}{K}.$
Here $H(\pi)=-\sum_{s}\pi_{s}\ln\pi_{s}$ is the Shannon entropy of the quantized source.
Moreover, we can decompose $H(\pi)$ as
$H(\pi)=H(p)-\sum_{j=1}^{m}\pi_{A_{j}}\ln v_{j}\;+\;D_{\mathrm{KL}}(\rho\|\bar{\rho}),$
where $v_{j}=\int_{A_{j}}dx$ and $\bar{\rho}=\sum_{j}(\pi_{A_{j}}/v_{j})\mathbf{1}_{A_{j}}$.
Combining, we finally obtain
$\bar{\ell}=\big(H(p)-\sum_{j=1}^{m}\pi_{A_{j}}\ln v_{j}\big)\;+\;D_{\mathrm{KL}}(\rho\|\bar{\rho})\;+\;D_{\mathrm{KL}}(\pi\|q)\;+\;\ln\frac{1}{K},$
which implies (by Gibbs’ inequality $D_{KL}\geq 0$ and Kraft’s $K\leq 1$) that
$\bar{\ell}\geq H(p)-\sum_{j=1}^{m}\pi_{A_{j}}\ln v_{j}$
The terms of the above expressions can be interpreted as follows:
  • i.
    The leading term $H(p)-\sum_{j}\pi_{A_{j}}\ln v_{j}$ is the “resolution-adjusted entropy.”
  • ii.
    The nonnegative correction terms reflect: within-cell mismatch $D_{\mathrm{KL}}(\rho\|\bar{\rho})$, Huffman tree mismatch $D_{\mathrm{KL}}(\pi\|q)$, and Kraft slack $\ln(1/K)$.
  • iii.
    Equality holds if $\rho$ is uniform on each cell, the code is complete ($K=1$), and the Kraft distribution $q$ coincides with $\pi$.
For a uniform distribution on a region $U\subset\mathbb{R}$ of measure $V$, the mixed entropy is
$H(p)=h(\rho)=\ln V,$
exactly mirroring Boltzmann’s relation $S=k_{B}\ln\Omega$, with $\Omega$ being the number (or volume) of accessible microstates. Then $e^{H(p)}$ is the perplexity in information theory and the multiplicity of microstates in statistical physics.
For a quantized code $Q=(U,\{S\}_{u\in U})$ define its resolution perplexity as
$V_{Q}=\exp\!\Big(\sum_{j}\pi_{A_{j}}\ln v_{j}\Big).$
This plays the role of a coarse-graining scale: the effective cell size in which the source is observed. In physics, such a scale is set by Planck’s constant $h^{d}$ when passing from classical to quantum phase space.
Since $\bar{\ell}$ is the average code length in nats, then $e^{\bar{\ell}}$ behaves like an effective number of distinguishable states at the chosen resolution. That is, $\bar{\ell}$ gives us the effective alphabet size, or the number of macrostates.
All of the above implies that the inequality
$\bar{\ell}\geq H(p)-\log V_{Q},$
or, in the multiplicative form,
$V_{Q}\cdot\exp(\bar{\ell})\geq V$
which is directly analogous to an uncertainty principle:
[…]
We can also write
$\bar{\ell}\geq\log\left(\frac{V}{V_{Q}}\right),$
which means that the average code length of the quantized code cannot be less than the information needed to specify one cell among the number of resolution-cells covering the accessible region.
Then, the optimal average code length $\bar{\ell}^{*}$ of a $D$-ary code subject to the Kraft–McMillan inequality, satisfies a relation similar to Boltzmann’s identity. Namely,
$\bar{\ell}^{*}=k_{D}\cdot\log W,\text{ with }W=\frac{V}{V_{Q}}\text{ and }k_{D}=\frac{1}{\log D},$
where $W$ is the reciprocal of the proportion of observable (or encodable) macrostates to the volume of entire microstate space, while $k_{D}$ plays the role of the “coding Boltzmann constant”.
3.5 Appendix: Shannon’s Entropy Lower Bound
Let $\mathcal{C}\subset[1,+\infty)$ be a code as defined throughout the manuscript, and let its course distribution satisfy
$p(dx)=\sum_{w\in W}\delta_{w}(dx)+\rho(x)dx,$
where $W\cap\mathrm{supp}\,\rho=\emptyset$.
A length function $\ell:\mathcal{C}\to[0,\infty]$ is admissible if
$\int e^{-\ell(x)}\,d\nu(x)\;\leq\;1.$
Then the entropy is
$H(p)=-\sum_{w\in W}P(w)\log P(w)\;-\;\int_{\mathcal{C}}\rho(x)\log\rho(x)\,dx.$
We assume that $\mathcal{C}$ satisfies the Kraft condition:
$K=\int e^{-\ell(x)}\,d\nu(x)\leq 1,$
and call
$q(x)=\frac{e^{-\ell(x)}}{K}$
the Kraft distribution.
The latter is equivalent to
$\ell(x)=-\ln q(x)-\ln K,$
and thus we obtain
$\mathbb{E}_{p}[\ell]$ $=\int_{\mathcal{C}}\ell(x)\,p(dx)=-\int_{\mathcal{C}}\ln q(x)\,p(dx)-\ln Z$
$=H(p)+D_{\mathrm{KL}}(p\|q)-\ln Z\;\geq\;H(p),$
since $D_{\mathrm{KL}}(p\|q)\geq 0$ (by Gibbs’ inequality) and $\ln Z\geq 0$ (by assumption).
Thus, we have an analogue of Shannon’s lower entropy bound
$\mathbb{E}_{p}[\ell]\ \geq\ H(p),$
with equality if and only if $K=1$ and $e^{-\ell}=\frac{dp}{d\nu}$ ($\nu$-a.e.).

### v4

@@
 Abstract
-In the present paper we give a principled derivation of Elias’ Omega code by combining a constrained variational formulation of prefix coding with a renormalization flow on codeword distributions.
-Starting from a Lagrangian that minimizes average code length under the Kraft–McMillan constraint, we show that the implied distribution is a fixed point of a coarse–graining map, yielding the canonical iterated logarithm length, up to an additive constant.
-This establishes completeness and asymptotic optimality, and connects universal integer coding with coarse–grained entropy, uncertainty–type bounds, and multiplicity relations familiar from statistical physics.
-The renormalization operator induces a discrete flow that converges to the Elias fixed point for any admissible initialization, up to a bounded error, offering a clean bridge between information–theoretic constraints and RG–style scale invariance.
+In the present paper we give a derivation of Elias’ Omega code from physics principles by combining a constrained variational formulation of prefix coding with a renormalization flow on codeword distributions.
+Starting from a Lagrangian that minimizes average codelength under the Kraft–McMillan constraint, we show that the implied distribution is a fixed point of a coarse–graining map, yielding the canonical iterated log-sum length, asymptotically up to an additive constant.
+This establishes completeness and asymptotic optimality, and connects universal integer coding with coarse–grained entropy, uncertainty–type bounds, and entropy relations familiar from statistical physics.
 1 Introduction
-Among several possible encoding procedures for integers, in his paper [1] Elias introduces the so–called $\omega$ encoding.
-This is a universal prefix code obtained by the simple iterative procedure below.
-Let $s=\text{bin}(x)$ be the binary string representation of a decimal number $x$, and let $x=\text{dec}(s)$ be the decimal value corresponding to a binary string $s$.
-Algorithm 1 Elias’ Omega: Encoding
+Universal integer codes are a classical topic at the interface of information theory and the theory of algorithms.
+Elias’ family of $\gamma$–, $\delta$–, and $\omega$–codes gives practical schemes whose expected lengths track entropy without prior knowledge of the source law – a property known as universality [1].
+The structure of Elias’ $\omega$ exhibits a recursive nature: the length of an integer depends on the size of the integer and on the sizes of the integers describing that size, and so on.
+It can also be interpreted as scale-invariance, and thus enables a renormalization approach to the probability distribution of codelengths.
+Let $s=\text{bin}(x)$ be the binary string representation of a positive decimal number $x$, and let $x=\text{dec}(s)$ be the decimal value corresponding to a binary string $s$, possibly with leading zeros.
+Then Elias’ $\omega$ is defined by the following two algorithms [2].
+Algorithm 1 Elias’ $\omega$: Encoding
 1:A positive integer $N\geq 1$
@@
 9:return $C$
-Algorithm 2 Elias’ Omega: Decoding
+Algorithm 2 Elias’ $\omega$: Decoding
 1:Encoded binary string $S$
@@
 10:return NN
-Since Shannon’s foundational work [2] and the Kraft–McMillan inequality [3, 4], a central question has been how to construct codes that approach entropy without knowledge of the true source distribution: such codes are called universal.
-Elias’s $\gamma$, $\delta$, and $\omega$ codes are all universal [1].
-In this note, we show that $\omega$ follows directly from a constrained Lagrangian optimization combined with a scale‐invariance requirement for the implied distribution, yielding a clean derivation from fundamental physical and code-theoretic principles.
+The code description given above implies the following curious identity about its codeword lengths.
+Lemma 1.1 (Elias $\omega$ codelength).
+Let $n_{0}=n$ and recursively define $n_{j+1}=\lfloor\log_{2}n_{j}\rfloor+1$. Stop at the first $m$ such that $n_{m}=1$. Then the Elias $\omega$ codelength is
+$\ell_{\omega}(n)\;=\;1+\sum_{j=0}^{m-1}\big(\lfloor\log_{2}n_{j}\rfloor+1\big).$ (1)
+Consequently,
+$\ell_{\omega}(n)\;=\;\sum_{j=0}^{m-1}\log_{2}n_{j}\;+\;O(\log^{*}n),$ (2)
+where $\log^{*}n$ is the iterated base-$2$ logarithm.
+The $\omega$–code is known to be prefix-free, meaning that no codeword can represent an initial segment of another codeword.
+Such codes are uniquely decodable since there is no ambiguity between the source elements encoded in a sequence and the sequence of codewords representing them.
+Being prefix-free turns out to be equivalent to the classical Kraft–McMillan [3, 4] conditions on the code word lengths.
+This condition turns out to be equivalent to a specific bound on the statistical sum with respect to a Gibbs measure.
+We recast Elias’ $\omega$–code in a language closer to statistical physics by combining
+  1. (i)
+    a Lagrangian formulation of optimal prefix coding under the Kraft–McMillan conditions [3, 4, 5], and
+  2. (ii)
+    a renormalization map implementing coarse-graining of the codeword distribution, and normalization.
+The fixed point of the induced flow on codelengths is exactly the $\omega$–codelength; moreover, this fixed point has a wide basin of attraction.
+The Lagrangian approach has been used before by Kafri to analyze Benford’s law [6], and by Visser [7] to derive Zipf’s law from the Maximum Entropy Principle [8, 9].
+The renormalization approach, to the best of our knowledge, has not appeared in the earlier literature.
+However, it is very naturally connected to the original definition of Elias’ $\omega$–code and its recursive nature.
+The encoding problem becomes a transparent toy model where scale-invariance is enforced by information-theoretical constraints (being uniquely decodable) rather than dynamics, linking two methodological pillars of statistical physics: variational principles and renormalization flows [10, 11, 12].
+Moreover, our analysis applies to codes with both discrete and continuous components of their codelength distribution.
+The usual information constraint for a discrete code that we use is the classical Kraft–McMillan inequality that guarantees the code is uniquely decodable.
+Reformulating it in terms of a Gibbs measure allows writing down the continuous analogue.
+Moreover, a continuous code can be quantized by binning together some of the codewords in order to achieve a series of continuous approximations with varying resolution.
+At each resolution level, the classical discrete theory applies, while This procedure recovers information-theoretical analogues of both classical Heisenberg’s (resolution–complexity tradeoff in code quantization) and Boltzmann’s (persistence of entropy at each resolution level) relations.
+In this work, we achieve several goals:
+  • give a variational derivation showing that the optimizer saturates the Kraft–McMillan inequality and identifies the implied codeword probability law – a property otherwise called completeness (Theorem 2.1);
+  • quantify asymptotic optimality $\mathbb{E}[\ell]\leq H(p)+O(1)$, with Shannon’s lower bound $\mathbb{E}[\ell]\geq H(p)$, and interpret the discrepancies using physics analogies: code quantization, Heisenberg’s and Boltzmann’s relations (Theorem 2.2);
+  • define a simple renormalization operator on codeword probability laws and the induced map on lengths; we prove that its fixed point is the iterated-log form of $\omega$ and that the flow is attractive up to a reasonably small error (Theorem 2.3).
 2 Optimal codes from Lagrangians and Renormalization
@@
 $p(dx)\;=\;\sum_{w\in W}P(w)\,\delta_{w}(dx)\;+\;\rho(x)dx,$ (3)
-i.e. a weighted sum of delta-functions together with a continuous probability density. Here $W$ is at most countable, $\rho\geq 0$ is integrable, and
+i.e. a weighted sum of delta-functions together with a continuous probability density. Here $W$ is at most countable, $\rho\geq 0$ is integrable, $W\cap\mathrm{supp}\,\rho=\emptyset$, and
 $\sum_{w\in W}P(w)+\int_{\mathcal{C}}\rho(x)\,dx=1.$ (4)
-Moreover, $W\cap\mathrm{supp}\,\rho=\emptyset$.
 Fix an output alphabet $\Sigma_{D}$ of size $D\geq 2$, and let $\Sigma_{D}^{*}$ be its Kleene closure.
@@
 A prefix code (or a prefix-free code) is a countable set $U\subset\Sigma_{D}^{*}$ of finite strings over a $D$-ary alphabet such that no $u\in U$ is a prefix of another $v\in U$.
-Let us formally recast the Kraft–McMillan inequality [3, 4] for prefix codes in the continuous form:
+Let us recast the Kraft–McMillan inequality [3, 4] for prefix codes in the continuous form:
 $\int_{\mathcal{C}}\exp(-\ell(x))\nu(dx)\leq 1,$ (6)
 where
-$\nu(dx)=\sum_{w\in W}\delta_{w}(dx)+dx,$ (5)
+$\nu(dx)=\sum_{w\in W}\delta_{w}(dx)+dx$ (7)
 is the Kraft measure.
@@
 $\mathcal{L}(\ell)=\int_{\mathcal{C}}\ell(x)p(dx)-\lambda\left(1-\int_{\mathcal{C}}\exp(-\ell(x))\nu(dx)\right)\to\min_{\ell},$ (8)
-with $\lambda>0$. This implies that violating the Kraft–McMillan inequality would imply a penalty on the Lagrangian.
+with $\lambda>0$. This implies that violating the Kraft–McMillan inequality would impose a penalty on the Lagrangian.
 The whole Lagrangian will then minimize the average codelength under the prefix-free constraint. This entropy–optimality interpretation is classical [8, 9, 5].
@@
 while
-$1=p(\mathcal{C})=\int_{\mathcal{C}}r(x)\nu(dx)=\lambda\int_{\mathcal{C}}\exp(\ell(x))\nu(dx)=\lambda.$ (9)
+$1=p(\mathcal{C})=\int_{\mathcal{C}}r(x)\nu(dx)=\lambda\int_{\mathcal{C}}\exp(-\ell(x))\nu(dx)=\lambda.$ (11)
 Indeed, the complementary slackness condition implies that the code so derived has to saturate Kraft’s inequality, and thus be complete (see [14] for fundamentals of convex analysis).
@@
 However, we shall see that universality will readily follow from this simple distributional condition.
-The process used below is in essence the one of “coarse‐graining” by replacing $x$ with $\log x$ followed by “renormalizing” the resulting probability (or code length).
+The process used below is in essence the one of coarse‐graining by replacing $x$ with $\log x$ followed by renormalizing the resulting probability (or codelength).
 However, it appears to be technically much simpler than the classical renormalization techniques in physics [10, 11, 12] due to its code-theoretical nature.
-Thus, put
+Let us put
 $\mathcal{R}p(\exp(x))=p(x)\cdot\exp(-x),$ (13)
@@
 $\ell(\exp(x))=\ell(x)+x.$ (15)
-The above, together with the natural constraint $\ell(x)\geq 0$, implies that $\ell(x)$ is given by the sum of iterated logarithms $\log(\log(\ldots\log(x)\ldots)$ that continues until we obtain a negative number.
+The above, together with the natural constraint $\ell(x)\geq 0$, implies that $\ell(x)$ is given by the sum of nested logarithms $\log(\log(\ldots\log(x)\ldots)$ that continues until we obtain a negative number.
 This implies that
@@
 Then the codelength of $\mathcal{C}$ behaves as $\ell(x)=\log x+o(\log x)$, for $x\gg 1$, and $\mathcal{C}$ is asymptotically optimal: $H(p)\leq\mathbb{E}[\ell]\leq H(p)+O(1)$.
-We shall call any code satisfying Theorems 2.1 - 2.2 an Elias’ Omega code, for short.
+We shall call any code satisfying Theorems 2.1 - 2.2 an $\omega$–type code, for short.
 2.3 Renormalization Flow Universality
@@
 $\mathcal{R}[p](\exp(x))\;=\;p(x)\,\exp(-x),$ (20)
-where the map $\log:\exp(x)\to x$ implements a coarse‐graining step or scale reduction, while the factor $\exp(-x)$ ensures $\int_{0}^{\infty}\mathcal{R}[p](x)\,dx=1$.
+where the map $\log:\exp(x)\to x$ implements a coarse‐graining step or scale reduction, while the factor $\exp(-x)$ ensures
+$\int_{1}^{\infty}\mathcal{R}[p](x)\,dx=\int^{\infty}_{0}\mathcal{R}[p](\exp(x))=$
+$=\int^{\infty}_{0}p(x)\exp(-x)\cdot\exp(x)dx=$
+$=\int^{1}_{0}0\cdot dx+\int^{\infty}_{1}p(x)dx=1,$
+where we assume $p(x)\equiv 0$ outside of the usual codeword range $1,+\infty)$.
 Assuming that $p(x)$ is implied, we have the induced operator for the codelength $\ell(x)$ is
-$T[\ell](x)\;=\;\ell\bigl(\log x\bigr)\;+\;\log x.$ (19)
+$\mathcal{T}[\ell](x)\;=\;\ell\bigl(\log x\bigr)\;+\;\log x.$ (21)
 Starting from any given codelength function $\ell_{0}(x)$, we can transform it by setting
-$\ell_{k+1}(x)=T[\ell_{k}](x)=\ell_{k}(\log x)+\log x,$ (20)
+$\ell_{k+1}(x)=\mathcal{T}[\ell_{k}](x)=\ell_{k}(\log x)+\log x,$ (22)
 for $k=0,1,2,\ldots$, which gives us a discrete flow.
-The fixed point $\ell^{*}(x)$ of this flow is exactly the Elias’ code length discussed above:
-$\ell_{*}(x)=T[\ell_{*}](x)=\ell_{*}(\log x)+\log x.$ (21)
-Recursively unrolling this relation yields exactly
+The fixed point $\ell^{*}(x)$ of this flow is almost the Elias’ codelength from Lemma 1.1. Namely, we have
+$\ell_{*}(x)=\mathcal{T}[\ell_{*}](x)=\ell_{*}(\log x)+\log x,$ (23)
+and recursively unrolling this relation yields
 $\ell_{*}(x)=\log x\;+\;\log(\log x)\;+\;\log\bigl(\log(\log x)\bigr)\;+\;\cdots,$ (24)
-terminating when the next $\log$ argument drops below $1$, i.e. the “0” termination of Elias’ Omega encoding.
+terminating when the next $\log$ argument drops below $1$, which is directly analogous to the “0” termination of Elias’ $\omega$ encoding.
+If we rewrite Elias’ $\omega$ codelength using natural logarithms instead of base-$2$ logarithms111i.e. we measure codelengths in nats rather than in bits., we get
+$\ell_{\omega}(x)=\ell_{*}(x)+O(\log^{*}x)=\ell_{*}(x)+o(\log(x)^{\varepsilon}),$ (25)
+where $\log^{*}(x)$ is the iterated logarithm of $x$ known for its very slow growth that guarantees the above equality for any $\varepsilon>0$.
 Since $\log x\leq x/2$ for any $x\geq 1$, we get that
 $\ell_{*}(x)\leq\log x\cdot\sum^{m}_{i=0}\frac{1}{2^{i}}\leq 2\log x$ (26)
-where $m=\log^{*}x$, the iterated logarithm of $x$, known for its very low rate of growth. This implies that
-$\ell_{*}(x)=O(\log x),$ (24)
-all the while being a self-delimiting code unlike the usual binary encoding, which obviously also has $O(\log x)$ code length.
+with $m=\log^{*}x$. This, together with the obvious $\ell_{*}(x)\geq\log x$, implies that
+$\ell_{*}(x)=\Theta(\log x),$ (27)
+all the while being a self-delimiting code unlike the usual binary encoding, which obviously also has $\Theta(\log x)$ codelength.
 Moreover, given a starting non-negative function $\ell_{0}(x)$ on the codewords of $\mathcal{C}$, repeated application of the map
-$T:\ell\mapsto\ell\circ\log+\log$ (25)
+$\mathcal{T}:\ell\mapsto\ell\circ\log+\log$ (28)
 flows to the point
@@
 Thus
-$|\ell_{*}(x)-\ell_{m}(x)|\leq\ell_{0}(x_{m})=O(1),$ (27)
+$|\ell_{*}(x)-\ell_{m}(x)|\leq\ell_{0}(x_{m})=\text{const},$ (30)
 since $\ell_{0}(x_{m})$ is simply the codelength of the termination marker.
-We can summarize our findings as the following propositions.
-Proposition 2.3.
-Let $\mathcal{C}$ be an Elias’ Omega code with code length $\ell(x)$ for $x\in\mathcal{C}$. Then
-$\ell(x)=\log x+\log^{(2)}x+\ldots+\log^{(m)}x,$
-with $m=\log^{*}x$, the iterated logarithm of $x$.
-Proposition 2.4.
-Let $\ell_{0}(x)$ be some non-negative function on the codewords of an Elias’ Omega code $\mathcal{C}$.
+We can summarize our findings as the following theorem.
+Theorem 2.3.
+Let $\ell_{0}(x)$ be some non-negative function on the codewords of an $\omega$–type code $\mathcal{C}$.
 Then the repeated iterations of the renormalization map $\ell\mapsto\ell\circ\log+\log$ almost converge to the canonical codelength.
-Namely, if $m=\log^{*}x$, let $\ell_{m}=T^{(m)}[\ell]$, and then
-$|\ell(x)-\ell_{m}(x)|<O(1)$
-independent of $x$.
+Namely, let $m=\log^{*}x$ and $\ell_{m}=\mathcal{T}^{(m)}[\ell]$.
+Then
+$|\ell_{w}(x)-\ell_{m}(x)|=\mathrm{const}+o(\ell_{\omega}(x)^{\varepsilon}),$
+for any $\varepsilon>0$.
 2.4 Passing to the Discrete Encoding
-Naturally, if we now recast our observation into the discrete setting, Elias’ Omega encoding [1] will provide a practical realization with very close properties.
-As it was mentioned before, Elias’ Omega encoding has to be complete.
-Indeed, let $L(n)$ be the length of $n$ in Elias’ Omega encoding, and let $\beta(n)=\lfloor\log_{2}n\rfloor+1$ be the binary length of $n$.
+Naturally, if we now recast our observation into the discrete setting, Elias’ $\omega$ encoding [2] will provide a practical realization with very close properties.
+As it was mentioned before, Elias’ $\omega$ encoding has to be complete.
+Indeed, let $\ell_{\omega}(n)$ be the length of $n$ in Elias’ $\omega$ code, and let $\beta(n)=\lfloor\log_{2}n\rfloor+1$ be the binary length of $n$.
 Then we have
-$L(n)=L(\beta(n)-1)+\beta(n).$ (28)
+$\ell_{\omega}(n)=\ell_{\omega}(\beta(n)-1)+\beta(n).$ (31)
 Let also
@@
 We have that $I_{k}$ contains $2^{k-1}$ elements, and thus
-$S_{k}=\sum_{n\in I_{k}}2^{-L(n)}=\sum_{n\in I_{k}}2^{-L(k-1)-k}=2^{-L(k-1)-1}.$ (30)
-Then for the Kraft’s sum $S=\sum_{n\in\mathbb{N}}2^{-L(n)}$ we get
-$S=\sum^{\infty}_{k=1}S_{k}=S_{1}+\sum^{\infty}_{k=2}S_{k}=\frac{1}{2}+\sum^{\infty}_{k=2}2^{-L(k-1)-1}=\frac{1}{2}+\frac{1}{2}\sum^{\infty}_{k=1}2^{-L(k)}=\frac{1+S}{2}.$ (31)
+$S_{k}=\sum_{n\in I_{k}}2^{-\ell_{\omega}(n)}=\sum_{n\in I_{k}}2^{-\ell_{\omega}(k-1)-k}=2^{-\ell_{\omega}(k-1)-1}.$ (33)
+Then for the Kraft’s sum $S=\sum_{n\in\mathbb{N}}2^{-\ell_{\omega}(n)}$ we get
+$S=\sum^{\infty}_{k=1}S_{k}=S_{1}+\sum^{\infty}_{k=2}S_{k}=\frac{1}{2}+\sum^{\infty}_{k=2}2^{-\ell_{\omega}(k-1)-1}=\frac{1}{2}+\frac{1}{2}\sum^{\infty}_{k=1}2^{-\ell_{\omega}(k)}=\frac{1+S}{2}.$ (34)
 This readily implies that $S=1$. The real rate of convergence is very slow though: in our computations we have reached (see [17])
-$\sum^{2^{2^{24}}}_{n=1}2^{-L(n)}\approx 0.9697265625\ldots$ (32)
+$\sum^{2^{2^{24}}}_{n=1}2^{-\ell_{\omega}(n)}\approx 0.9697265625\ldots$ (35)
 Acknowledgments
 This material is based upon work supported by the Google Cloud Research Award number GCP19980904.
-3 Appendix: Kraft–McMillan and Quantization
+3 Appendix: Continuous codes and Quantization
 3.1 Mixed Source Distributions
 Let $p(x)$ be a probability measure on the set of codewords $\mathcal{C}=[1,+\infty)$ of the form
-$p(x)\;=\;\sum_{w\in W}P(w)\,\delta_{w}(x)\;+\;\rho(x),$
+$p(dx)\;=\;\sum_{w\in W}P(w)\,\delta_{w}(dx)\;+\;\rho(x)dx,$
 where $W$ is at most countable, $\rho\geq 0$ is integrable, $W\cap\mathrm{supp}\,\rho=\emptyset$, and
@@
 A prefix code (or a prefix-free code) is a countable set $U\subset\Sigma_{D}^{*}$ of finite strings over a $D$-ary alphabet such that no $u\in U$ is a prefix of another $v\in U$.
-Instead of the $D$-ary code length $\ell_{D}(x)$ for $x\in\mathcal{C}$, we shall use the baseless (more precisely, base–invariant) code length $\ell(x)=\ell_{D}(x)/\log D$.
+Instead of the $D$-ary codelength $\ell_{D}(x)$ for $x\in\mathcal{C}$, we shall use the baseless (more precisely, base-invariant) codelength $\ell(x)=\ell_{D}(x)/\log D$.
 In the mixed (discrete and continuous) distribution case, we shall assign source outcomes to these finite strings by grouping outcomes into finitely many coding cells represented either by singleton atoms or quantization cells of the density.
@@
 Thus, there exists a quantized prefix code for the mixed-distribution code $\mathcal{C}$, for every admissible choice of the index set $S$ as above.
-3.4 Heisenberg–style and Boltzmann–style relations
+3.4 Heisenberg-style and Boltzmann-style relations
 Let the Kraft sum be $K=\sum_{s}e^{-\ell_{s}}$, and let us define the Kraft distribution as
@@
 $\bar{\ell}=\sum_{s\in S}\pi_{s}\ell_{s}\;=\;H(\pi)\;+\;D_{\mathrm{KL}}(\pi\|q)\;+\;\ln\frac{1}{K}.$
-Here $H(\pi)=-\sum_{s}\pi_{s}\ln\pi_{s}$ is the Shannon entropy of the quantized source.
+Here
+$H(\pi)=-\sum_{s}\pi_{s}\ln\pi_{s}$
+is the Shannon entropy of the quantized source, and
+$D_{\mathrm{KL}}(\pi\|q)=\sum_{s}\pi_{s}\log\frac{\pi_{s}}{q_{s}}$
+is the Kullback–Leibler (KL) divergence of two distributions [5].
 Moreover, we can decompose $H(\pi)$ as
@@
 3.5 Appendix: Shannon’s Entropy Lower Bound
-Let $\mathcal{C}\subset[1,+\infty)$ be a code as defined throughout the manuscript, and let its course distribution satisfy
+Let $\mathcal{C}\subset[1,+\infty)$ be a code as defined throughout the manuscript, and let its coarse distribution satisfy
 $p(dx)=\sum_{w\in W}\delta_{w}(dx)+\rho(x)dx,$
@@
 A length function $\ell:\mathcal{C}\to[0,\infty]$ is admissible if
-$\int e^{-\ell(x)}\,d\nu(x)\;\leq\;1.$
+$\int e^{-\ell(x)}\,\nu(dx)\;\leq\;1.$
 Then the entropy is
@@
 We assume that $\mathcal{C}$ satisfies the Kraft condition:
-$K=\int e^{-\ell(x)}\,d\nu(x)\leq 1,$
+$K=\int e^{-\ell(x)}\,\nu(dx)\leq 1,$
 and call